# DSPy Assertions: Computational Constraints for Self-Refining Language Model Pipelines

### 20 Dec 2023 v1

@@ Abstract
Chaining language model (LM) calls as composable modules is fueling a new powerful way of programming. However, ensuring that LMs adhere to important constraints remains a key challenge, one often addressed with heuristic “prompt engineering”. We introduce LM Assertions, a new programming construct for expressing computational constraints that LMs should satisfy. We integrate our constructs into the recent DSPy programming model for LMs, and present new strategies that allow DSPy to compile programs with arbitrary LM Assertions into systems that are more reliable and more accurate. In DSPy, LM Assertions can be integrated at compile time, via automatic prompt optimization, and/or at inference time, via automatic self-refinement and backtracking. We report on two early case studies for complex question answering (QA), in which the LM program must iteratively retrieve information in multiple hops and synthesize a long-form answer with citations. We find that LM Assertions improve not only compliance with imposed rules and guidelines but also enhance downstream task performance, delivering intrinsic and extrinsic gains up to $35.7\%$ and $13.3\%$, respectively. Our reference implementation of LM Assertions is integrated into DSPy at dspy.ai.
@@ 1 Introduction
Language models (LMs) now power various applications in Natural Language Processing (NLP) from conversational agents to writing assistants. However, the probabilistic nature of LMs often results in outputs that may not align with the constraints of the domain or the larger pipeline in which the LM is used. To address this, researchers have explored various techniques, including applying constrained decoding (Hokamp & Liu, 2017; Hu et al., 2019), monitoring model assertions (Kang et al., 2020), exploring approaches for self-reflection and tree search (Madaan et al., 2023; Shinn et al., 2023; Yao et al., 2023), or even building domain-specific languages and systems such as LMQL (Beurer-Kellner et al., 2023) and Guardrails (Rebedea et al., 2023) to steer LMs towards more controllable outputs.
@@
Recently, several LM application frameworks like LangChain (Chase, 2022) and LM programming models like DSPy (Khattab et al., 2022, 2023) have been proposed, providing developers with simpler interfaces to build complex LM pipelines. While these frameworks offer a number of key features to control LM outputs (e.g., DSPy’s automatic compiler), they lack the expressiveness to specify arbitrary computational constraints on LM pipelines and enable them to learn from these constraints and to introspectively self-refine outputs. While some of this may be achieved via pain-staking “prompt engineering” or other ad-hoc guidance strategies, this would be labor-intensive and in general conflates high-level program design with the low-level exploration of teaching LMs how to follow constraints and using them iteratively for self-refinement.
We introduce LM Assertions, a novel programming construct designed to enforce user-specified properties on LM outputs within a pipeline framework. Drawing inspiration from runtime assertions and program specifications in traditional programming, LM Assertions are expressed as boolean conditions that reflect the desired characteristics of LM outputs. Besides serving as conventional runtime monitors, LM Assertions create multiple new ways to improve LM programs. First, they can be used to facilitate runtime self-refinement in LM pipelines. When an LM Assertion fails, instead of simply terminating, the pipeline backtracks and retries the failing pipeline module. LM assertions provide feedback on retry attempts; they inject erring outputs and error messages to the prompt to introspectively self-refine outputs. Figure 1 illustrates this self-refinement process in a DSPy pipeline. Second, LM Assertions can enable guided prompt optimizers at compile time. Integrated with existing automatic prompt optimizers, like in DSPy, they can enable compiling harder few-shot examples, which tunes LM programs into systems that are more robust and accurate. In the same vein, the failing LM assertions can also be used to collect counter-examples for optimizers.
@@
We distinguish between two types of LM Assertions: (hard) Assertions and (soft) Suggestions, denoted by Assert and Suggest, respectively. Hard assertions represent critical conditions that, when violated after a maximum number of retries, cause the LM pipeline to halt, signaling a non-negotiable breach of requirements. On the other hand, suggestions denote desirable but non-essential properties; their violation triggers the self-refinement process, but exceeding a maximum number of retries does not halt the pipeline. Instead, the pipeline continues to execute the next module.
@@
We implement our work by extending the capabilities of the DSPy framework, a state-of-the-art framework for building and automatically optimizing declarative LM pipelines, by integrating LM Assertions as a module. This integration not only enables DSPy programs to self-refine and produce outputs that adhere to specific guidelines but also simplifies the debugging experience, providing developers with a clearer understanding of LM behavior within complex pipelines. In addition, by combining LM Assertions with prompt optimizers in DSPy, we envision the potential to bootstrap harder few-shot demonstrations to make the pipeline more robust and performant.
We evaluate our approach on the downstream multi-hop QA task of generating brief 1–5 word answers for questions within the HotPotQA dataset. We apply LM Assertions to refine retrieval queries by imposing requirements of conciseness and distinctness from past queries to ensure precise information retrieval and answer correctness. Our experiments show that by incorporating the assertion, search queries are $27.2\%$$35.7\%$ more likely to be distinct from previous queries; thus, retrieval queries can fetch more useful information in the context. As a result, the useful passages retrieved increase $4.2\%$$13.3\%$, and the final answer correctness increases $1.3\%$$1.4\%$.
We build on this with an enhancement to the task: generating citation-inclusive paragraphs to answer the questions (Gao et al., 2023). This task exemplifies the utility of LM Assertions, requiring the LM to adhere to strict formatting rules and maintain fidelity to cited sources. By incorporating computational constraints into the DSPy framework, we demonstrate significant improvements in the quality of generated content with more than $16.7\%$ more faithful citations, showcasing the potential of LM Assertions to elevate the performance of LMs in tasks that demand precision and adherence to domain-specific rules.
@@ 2 Background and Motivation
The goals of LM Assertions are general and can be applied to any LM program. Due to its modular paradigm, flexibility, and extensibility, we implement our work as extensions to the state-of-the-art DSPy (Khattab et al., 2023) framework. Below, we briefly describe the DSPy programming model for building declarative LM pipelines and compiling them into auto-optimized prompt (or finetune) chains. We then sketch a realistic, motivating example for LM Assertions and show their usefulness for self-refinement in LM pipelines.
@@ 2.1 The DSPy Programming model
First, instead of free-form string prompts, a programmer using DSPy will define a signature to declaratively specify what an LM needs to do. For instance, a module may need to consume a question and return an answer, as shown below:
@@
    rationale_field = dspy.OutputField(prefix="Reasoning: Think step by step.")
    signature = dspy.Signature(signature).prepend_output_field(rationale_field)
@@
DSPy modules can be composed in arbitrary pipelines by first declaring the modules needed at initialization and then expressing the pipeline with arbitrary code that calls the modules in a forward method (as shown in the ChainOfThought module above and the MultiHopQA program in Section 2.2). Finally, DSPy provides teleprompters, which are optimizers that can, among other things, automate generating good quality demonstrations (few-shot examples) or instructions for a task given a metric to optimize. We may also refer to the few-shot example selection process as compiling the LM pipeline application.
Challenges.
DSPy signatures provide type hints that softly shape LM’s behavior. However, the framework currently lacks constructs developers can use to specify arbitrary computational constraints the pipeline must satisfy. Additionally, one can imagine the LM pipeline using these constraints to refine its outputs and to teach the LM to respect these specifications at compile time.

### 02 Feb 2024 v2

@@ Abstract
Chaining language model (LM) calls as composable modules is fueling a new way of programming, but ensuring LMs adhere to important constraints requires heuristic “prompt engineering.” We introduce LM Assertions, a programming construct for expressing computational constraints that LMs should satisfy. We integrate our constructs into the recent DSPy programming model for LMs and present new strategies that allow DSPy to compile programs with LM Assertions into more reliable and accurate systems. We also propose strategies to use assertions at inference time for automatic self-refinement with LMs. We report on four diverse case studies for text generation and find that LM Assertions improve not only compliance with imposed rules but also downstream task performance, passing constraints up to $164\%$ more often and generating up to $37\%$ more higher-quality responses.
@@ 1 Introduction
Language models (LMs) now power various applications, from conversational agents to writing assistants. However, the probabilistic nature of LMs often results in outputs that may not align with the constraints of the domain or the larger pipeline in which the LM is used. To address this, researchers have explored various techniques, including applying constrained decoding (Hokamp & Liu, 2017; Hu et al., 2019), exploring approaches for self-reflection and tree search (Madaan et al., 2023; Shinn et al., 2023; Yao et al., 2023), building domain-specific languages and like LMQL (Beurer-Kellner et al., 2023), or monitoring models with assertions and guardrails (Kang et al., 2020; Rebedea et al., 2023) to steer LMs towards more controllable outputs.
@@
Recently, several LM frameworks like LangChain (Chase, 2022) and DSPy (Khattab et al., 2022, 2024) provide developers with interfaces to build complex LM pipelines. Some offer several features to control LM outputs, e.g., DSPy can optimize multi-stage prompts to maximize a target metric. However, they cannot take arbitrary computational constraints on LM pipelines to either teach the LM to follow such constraints and/or to introspectively self-refine outputs. While some of this may be achieved via pain-staking “prompt engineering” or other ad-hoc guidance strategies, such efforts are labor-intensive and conflate the high-level design of new AI systems with the low-level exploration of teaching LMs how to follow constraints.
We propose LM Assertions, a novel programming construct designed to enforce user-specified properties on LM outputs within a pipeline. Drawing inspiration from runtime assertions and program specifications in traditional programming, LM Assertions are expressed as boolean conditions that express the desired characteristics of LM outputs. We integrate LM Assertions into the recent DSPy programming model. Besides serving as conventional runtime monitors, LM Assertions enable multiple novel assertion-driven optimizations to improve LM programs.
Assertion-driven backtracking. LM Assertions can facilitate runtime self-refinement in LM pipelines at inference time. When a constraint fails, we allow the pipeline to backtrack and retry the failing module. LM Assertions provide feedback on retry attempts; they inject erring outputs and error messages to the prompt to introspectively self-refine outputs. Figure 1 illustrates this within a DSPy pipeline.
class MultiHopQAWithAssertions(dspy.Module):
  def forward(self, question):
    context, queries = [], [question]
    for hop in range(2):
      dspy.Suggest(len(query) < 100,
          "Query should be less than 100 characters")
       dspy.Suggest(is_query_distinct(query, queries),
          f"Query should be distinct from {queries}")
      context += self.retrieve(query).passages
      queries.append(query)
Figure 1: DSPy program with LM Assertions for multi-hop question-answering task with a retriever. We introduce two soft assertions (suggestions): (1) query to retriever should be less than 100 characters; (2) query to retriever should differ from previous queries. For instance, if the second suggestion fails, DSPy will construct a new prompt to retry the generate_query module with additional fields, highlighting the previously generated query and a user-defined error message to help the LM refine its generation.
Assertion-driven example bootstrapping. LM Assertions can enable guided prompt optimizers at compile time. Integrated with existing automatic prompt optimizers in DSPy, they can enable generating harder few-shot examples, which can teach LM programs to conduct challenging steps.
Counterexample bootstrapping. During prompt optimization and example bootstrapping, another important contribution of LM Assertions is to develop demonstrations that contain failed examples and traces to fix the errors. When the counterexamples are mixed with bootstrapped few-shot examples, the LM is more likely to avoid the same mistakes without assertion-driven backtracking.
@@
We propose two types of LM Assertions: (hard) Assertions and (soft) Suggestions, denoted by Assert and Suggest, respectively. Hard assertions represent critical conditions that, when violated after a maximum number of retries, cause the LM pipeline to halt, signaling a non-negotiable breach of requirements. On the other hand, suggestions denote desirable but non-essential properties; their violation triggers the self-refinement process, but exceeding a maximum number of retries does not halt the pipeline. Instead, the pipeline continues to execute the next module.
@@
We implement our work atop DSPy, a state-of-the-art framework for building and automatically optimizing declarative LM pipelines. The integration enables us to design and implement our three new assertion-driven optimizations. It enables DSPy programs to self-refine and produce outputs that adhere to specific guidelines. It simplifies debugging, providing developers a clearer understanding of LM behavior in complex pipelines. In addition, by combining LM Assertions with prompt optimizers in DSPy, we can bootstrap better few-shot examples and counterexamples to assertions to make the pipeline more robust and performant.
We evaluate the effectiveness of LM Assertions on four varied knowledge-intensive tasks: multihop question answering (MultiHopQA), long format question answering (LongFormQA), formatted quiz generation (QuizGen), and valid tweet generation (TweetGen). Our experiments show that LM Assertions and assertion-driven backtracking significantly improve LM programs from generating 37.6% well-formatted quizzes in JSON to 98.8% in QuizGen. Then, with assertion-driven example bootstrapping and counterexample bootstrapping, we see an increase from 2.0% to 73.0% in writing more engaging tweets in TweetGen and a consistent boost on almost all other assertion metrics. Finally, with LM Assertions and all assertion-driven optimizations, we see a maximum gain from 30.5% to 87.2% in generating valid quiz questions.
Our contributions are, first, introducing LM Assertions as an abstraction for language model programming. Then, we present three novel optimizations for LM pipelines: assertion-driven backtracking during inference to help models self-refine, assertion-driven example bootstrapping to generate and choose more effective few-shot examples, and counterexample bootstrapping to develop few-shot examples with erroneous results and corresponding fixes to help the model become more reliable at complying to developer-desired constraints.
@@ 2 Background and Motivation
The goals of LM Assertions are general and can be applied to any LM program. Due to its modular paradigm, flexibility, and extensibility, we implement our work as extensions to the state-of-the-art DSPy (Khattab et al., 2024) framework. Below, we briefly describe the DSPy programming model for building declarative LM pipelines and compiling them into auto-optimized prompt (or finetune) chains. We then sketch a realistic, motivating example for LM Assertions and show their usefulness for self-refinement in LM pipelines.
@@ 2.1 The DSPy Programming model
First, instead of free-form string prompts, a DSPy programmer will define a signature to specify what an LM needs to do declaratively. For instance, a module may need to consume a question and return an answer, as shown below:
@@
    rationale_field = dspy.OutputField(
        prefix="Reasoning: Think step by step.")
    signature = dspy.Signature(signature).
        prepend_output_field(rationale_field)
@@
DSPy modules can be composed in arbitrary pipelines by first declaring the modules needed at initialization and then expressing the pipeline with arbitrary code that calls the modules in a forward method (as shown in the ChainOfThought module above and the MultiHopQA program in Section 2.2). Finally, DSPy provides optimizers that automates generating good quality demonstrations (few-shot examples) or instructions for a task given a metric to optimize. We may also refer to the few-shot example selection process as compiling the LM pipeline application.
Challenges. DSPy signatures provide type hints that softly shape LM’s behavior. However, the framework currently lacks constructs developers can use to specify arbitrary computational constraints the pipeline must satisfy. Additionally, one can imagine the LM pipeline using these constraints to refine its outputs and to teach the LM to respect these specifications at compile time.